# supSupplementary References

### v1

@@ 3.3.1 Evaluating LLM Annotations
3.4 Task 3: Risk of Bias Support Judgment Selection
The previous task evaluates a model’s ability to extract information from biomedical papers to in order to justify assessments of bias. However, this approach alone is insufficient for a risk-of-bias assessment. The evaluation must also include the model’s ability to reason about the retrieved information. For example, a paper might report the loss to follow-up yet fail to provide reasons for missing data, which raises the possibility of attrition bias. Similarly, while a paper may state that it implemented double-blinding, expert reviewers could consider such blinding infeasible due to the nature of the trial, potentially introducing performance bias.
To address this, we propose a multiple-choice task where the model selects the correct support judgment. This task includes three synthetically generated hard options, three options derived from other papers’ support judgments for the same bias category, and the Cochrane reviewer’s actual judgment. We prompt GPT-4 to generate the three synthetic options by imitating support judgments from other papers concerning the same type of bias. These options are tailored to be paper-specific while maintaining the underlying reasoning. See Appendix D for details on construction of these synthetic judgments.
In order to ensure that all incorrect options were actually incorrect, we performed filtering in two stages. GPT-4 was used to perform preliminary filtering (the prompt is shown in Figure 13 of Appendix D). In the second stage, we manually reviewed all of the remaining options, and removed any that were not incorrect.
Figure 4: Risk of Bias Support Judgment Selection. In Task 3, the model is shown a paper, and asked which of 7 support judgments provides the best explanation of the paper’s risk of bias. (Only four options are shown here for illustration purposes.)
3.5 Task 4: Risk Level Determination
The final task directly evaluates a model’s ability to assess the risk of bias in biomedical research papers. It requires the model to categorize the risk level of a specific bias as high, low, or unclear. The input to the model includes the whole paper, the PICO of the study, the objective of the meta-analysis, and the definition of the bias, as illustrated in Figure 5.
The review authors use bias with definitions from Cochrane Handbook 5.1 Higgins and Green [2011], Cochrane Effective Practice and Organisation of Care Practice and of Care [EPOC], A Cochrane Risk of Bias Assessment Tool Sterne et al. [2014], or specific biases that are defined within the meta-analysis. For the last category, we extract the bias definition from the specific meta-analysis. (Bias definitions are also extracted for Tasks 2 and 3.)
This assessment requires synthesizing various elements identified in the previous tasks, including the accurate retrieval and interpretation of relevant sentences and the ability to critically evaluate the sufficiency and reliability of the evidence presented regarding each bias.
Figure 5: Risk Level Determination. In Task 4, the goal is to provide an assessment of the paper’s risk of specific biases.
@@ 4 Experiments
4.1 Experimental Models and Procedures
We evaluate GPT-4o OpenAI [2024], Claude3-Opus Anthropic [2024a], Claude3.5-Sonnet Anthropic [2024b], and Gemini-1.5-Pro Google [2024]. Due to the long-context nature of RoBBR’s tasks, with many datapoints exceeding 16k tokens, many open-sourced LLMs AI [2024], Poli et al. [2023], Li et al. [2023] are considered but none of which is equipped to evaluate RoBBR.
We use chain-of-thought as prompting strategy to stabilize the model performance. We optimize our prompts on a separate development set. See Appendix E for details on prompt optimization.
@@ 4.2 Evaluation metrics
For Task 1, study inclusion/exclusion, we report macro-F1, accuracy, and weighted accuracy.
For Task 2, Risk of Bias Sentence Retrieval, we report aspect-level recall metrics, which measure the proportion of the bias aspects that are covered by the retrieved sentences. Specifically, Bias Recall (BR) @ 3 measures the proportion aspects covered when three sentences are retrieved. BR @ Optimal measures recall when the “optimal" number of sentences are retrieved. The optimal number of sentences is defined as the fewest which are sufficient to cover all bias aspects.
For Task 3, Risk of Bias Support Judgment Selection, we report accuracy.
For Task 4, Risk Level Determination, we report accuracy and weighted accuracy.
4.3 Categorization
For Tasks 2, 3, and 4, we consider 8 categories of biases: selection bias, attrition bias, performance bias, detection bias, reporting bias, analysis bias, special bias (the biases that are meta-analysis specific), and other bias (i.e. biases that do not fall into any of the other categories). In order to reduce evaluation costs, we randomly sampled 50 task instances for some categories. It is important to note that systematic reviews vary in which biases they report. When the authors judge that a bias is not important for the objective of a review, they may decide not to evaluate a specific bias. As a result, not all categories have the same number of examples, and some categories have fewer than 50 examples. See Table 10 for detailed statistics.
5 Results
Table 2: Results for Task 1: Study Inclusion/Exclusion
F1 Accuracy Weighted Accuracy
Claude3-Opus 75.5 77.1 74.9
Claude3-Sonnet 80.6 81.1 80.8
Gemini-1.5-Pro 72.2 72.7 73.0
GPT-4o 75.5 75.6 77.6
[…]
Bias Type
Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
n = 337 n = 117 n = 57 n= 50 n = 50 n = 35 n = 13 n = 7 n = 23
Bias Recall@Optimal
Claude3-Opus 39.0 55.4 22.0 57.3 46.4 11.0 47.9 13.9 4.4
Gemini-1.5-Pro 41.1 57.2 28.7 49.1 48.8 23.3 54.2 13.9 3.3
GPT-4o 44.8 60.3 28.6 64.6 48.7 32.9 28.7 28.0 2.3
Bias Recall@3
Claude3-Opus 52.3 67.9 40.2 63.1 55.9 25.5 55.9 64.6 19.7
Gemini-1.5-Pro 56.1 69.7 44.1 73.7 65.3 35.6 56.3 35.5 6.7
GPT-4o 61.8 72.2 47.9 82.0 78.5 43.2 67.9 50.4 10.3
5.1 Performance on the Four Tasks
In Task 1 (Study Inclusion/Exclusion), shown in Table 2, Claude3-Opus and GPT-4o perform best in terms of F1 score. Claude3-Opus has higher accuracy and GPT-4o higher weighted accuracy.
In Task 2 (Sentence Retrieval Evaluation Results) 3, GPT-4o has stronger results than both Gemini-1.5 and Claude3-Opus for almost all bias types. However, in Task 3 (Support Judgment Selection Evaluation Results) 4, Claude3-Opus has the strongest results.
Tasks 2 and 3 are critical for evaluating the different capabilities of the LLMs. Task 2 is a long-context task, requiring retrieval of a non-redundant set of sentences from a 200-sentence paper. Task 3 requires complex reasoning to evaluate the retrieved sentences and identify the implications for risk of bias. Our results demonstrate that GPT-4o performs better in long-context retrieval tasks, while Claude3-Opus exhibits stronger reasoning abilities.
Figure 4 illustrates Claude3-Opus’s stronger reasoning skills. Options C and D present different reasoning based on the same information from the paper. Notably, GPT-4o is misled by a deceptive option, whereas Claude3-Opus correctly handles the multiple-choice question.
Task 4 (Risk Level Determination Evaluation Results) 5 provides an overall assessment of each model’s ability to determine the risk level, integrating all information from the previous tasks. The three models show comparable performance in this task.
[…]
Bias Type
Model Full Selection Attrition Performance Detection Reporting Analysis Special
n = 310 n = 86 n = 60 n = 52 n = 56 n = 50 n = 8 n = 6
Accuracy
Claude3-Opus 51.2 62.6 56.7 48.0 46.5 44.2 50.0 0.0
Claude3.5-Sonnet 60.2 69.8 79.9 55.8 44.8 51.8 62.5 0.0
Gemini-1.5-Pro 49.3 57.1 59.9 46.4 42.8 40.2 49.3 16.6
GPT-4o 45.8 52.3 63.1 42.5 37.7 28.0 74.5 0.0
[…]
Bias Type
Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
n = 382 n = 93 n = 50 n = 50 n = 50 n = 50 n = 18 n = 26 n = 50
F1
Claude3-Opus 47.7 54.3 34.1 46.7 37.4 23.0 69.9 27.9 27.9
Claude3.5-Sonnet 49.7 62.7 23.9 49.9 46.7 30.1 48.3 28.7 33.8
Gemini-1.5-Pro 43.1 53.9 26.0 43.7 46.8 30.4 44.6 20.1 22.8
GPT-4o 51.9 59.0 36.5 48.9 36.0 27.7 53.8 34.7 29.8
Accuracy
Claude3-Opus 56.8 64.6 58.1 48.2 41.7 40.0 88.9 53.8 72.2
Claude3.5-Sonnet 62.3 64.4 52.0 52.1 55.9 64.3 83.0 69.5 72.1
Gemini-1.5-Pro 49.3 62.5 42.4 46.5 48.1 56.0 77.5 42.4 23.7
GPT-4o 62.6 71.0 68.3 54.4 40.1 57.6 88.7 65.4 62.1
Weighted Accuracy
Claude3-Opus 49.5 55.3 36.2 48.0 46.5 30.4 69.3 28.2 32.4
Claude3.5-Sonnet 50.3 63.1 24.1 51.6 48.2 32.1 46.8 36.1 36.1
Gemini-1.5-Pro 47.9 56.3 34.0 44.4 54.2 36.4 43.6 22.2 32.5
GPT-4o 53.6 60.4 36.3 51.3 40.7 30.0 50.0 64.3 32.6
5.2 Performance across Bias Types
The three models exhibit significant variations in performance across different biases due to the varying challenges associated with each bias type. All models have stronger performance at assessing selection bias, as biomedical papers typically detail their randomization processes explicitly.
In contrast, assessing reporting bias requires a comprehensive understanding of the entire paper to determine whether all outcomes specified in the study protocol are reported. Evaluating attrition bias involves extracting data on participant dropout during the trial and assessing how this dropout impacts the study’s statistical estimates. The most challenging bias type to assess is "other bias," where the model must grasp all potential biases, such as those stemming from funding sources or bias due to lack of power. Successfully identifying them requires a thorough understanding of all possible biases a study might have, thereby necessitating the treatment of each bias type as a distinct task.
6 Related Work
Scientific and Biomedical Information Retrieval has been the subject of a significant body of work in recent years Jin et al. [2023], MacAvaney et al. [2020], Wang et al. [2023]. A subset of this work has been focused on benchmarks Ben Abacha and Demner-Fushman [2019], Krithara et al. [2023], Voorhees et al. [2021], Cohen et al. [2017].
Study Screening Many works have focused on the task of screening studies for inclusion in systematic reviews Cohen et al. [2006], Khabsa et al. [2016], Kontonatsios et al. [2017], El-Gayar et al. [2015], with more recent work often focusing on LLMs Wilkins [2023], Tran et al. , Guo et al. [2024], Na et al. [2024], Cai et al. [2023], Ye et al. [2024], Wang et al. [2024]. Huotala et al. [2024] conduct an evaluation of LLM paper screening, finding that few-shot chain-of-thought prediction performs comparably to human annotators. Robinson et al. [2023] leverage an instruction-tuned Guanaco-7B model on a large dataset of systematic reviews to exceed the screening performance of ChatGPT. Cohen et al. [2006] produce a dataset of screening judgments for 15 systematic drug reviews.
Study Risk of Bias Risk of bias identification requires a thorough reading of an entire paper. Marshall et al. [2015, 2016] train an SVM model with documents annotated for risk of bias, creating RobotReviewer system to predict risk of bias levels. Wang et al. [2022b] utilized a BERT-based pipeline for risk of bias prediction. However, several other works find lower agreement rates between LLMs and human annotators Lai et al. [2024], Hasan et al. [2024], Barsby et al. [2024].
Other LLM Extraction Efforts Recent years have seen LLMs benchmarked in extracting "ground truth" from systematic reviews, particularly in the context of data extraction Sun et al. [2024], Gartlehner et al. [2023], multi-document summaries Wang et al. [2022a], Wallace et al. [2020] and relevance prediction Al-Hussaini et al. [2022]. Another use case has been the extraction of PICOs (Participants, Interventions, Comparisons, Outcomes) from studies. While earlier methods achieved strong results Zafar et al. [2023], Hu et al. [2023], Wallace et al. [2016], Dhrangadhariya and Müller [2023], Jin and Szolovits [2020], Kang et al. [2019], a body of recent work has focused on extracting PICOs with LLMs. Fine-tuning Wadhwa et al. [2023] and prompt tuning Tang et al. [2024] have been used to enhance extraction.
Evaluation of Zero-shot LLM Extraction Kartchner et al. [2023] show that ChatGPT performed inconsistently on various meta-analysis extraction tasks. For instance, while it correctly extracted cancer type roughly 90% of the time, treatment type accuracy fell below 25%. This inconsistency is a considerable motive to more comprehensively understand LLM reliability on various meta-analysis extraction tasks.
Limitation Tasks in Scientific Documents Early efforts in assessing papers’ limitations focused on performing summary extraction and evaluating the result using human evaluation Cohan et al. [2018, 2022], Liu and Shah [2023].
For understanding scientific documents, identifying and retrieving a paper’s limitation is essential to understand the utility and viability of claims and evidences presented in the paper. Faizullah et al. [2024] evaluate fine-tuned embedding models and LLMs’ ability to generate limitation using the extracted limitation from the original paper.
Marshall et al. [2016] use direct quotes in risk of bias judgment to obtain annotations for review authors’ judgments, where exact matching texts are labeled as relevant to a risk of bias. Suster et al. [2023] present a quality assessment task to assess a biomedical paper using the Grading of Recommendation, Assessment, Development, and Evaluation with justifications.

### v2

@@ 4.3 Automatic Annotation Pipeline
Metrics Human & Human Human & Embedding p-value
Exact Accuracy 99.4 $\pm$ 0.2 98.5 $\pm$ 0.2 <10e3<\!10e{-3}
F1 Binary 73.4 $\pm$ 6.3 30.5 $\pm$ 8.1 <10e3<\!10e{-3}
Cohen’s $\kappa$ 73.1 $\pm$ 6.3 29.7 $\pm$ 8.2 <10e3<\!10e{-3}
Spearman’s $\rho$ 73.8 $\pm$ 6.0 29.8 $\pm$ 8.3 <10e3<\!10e{-3}
Table 2: Inter-annotator agreement. Top: comparison between two human teams and GPT-4. Bottom: comparison between two human teams and OpenAI-v3-large as a post-hoc analysis.
4.3.1 Evaluating Annotation Quality
We hypothesize a sentence matches an aspect is straightforward enough for GPT-4 to achieve human-level performance with a specialized prompt. Following the protocol introduced by Wang et al. (2023), we optimized the instruction prompt for GPT-4 on a development set (see Appendix D). We randomly sampled 50 papers, with each assigned a single aspect. Across all 50 papers, this resulted in a total of 13,575 (aspect, sentence) pairs. Four graduate student were tasked with annotating these pairs. The annotators were divided into two teams, with each team consisting of two annotators. The annotators’ training for annotation is described in Appendix B.1. Each annotator performed the annotation tasks individually given a pre-defined annotation guideline in Appendix B. Annotators from the same team then collaborated to resolve differences and eliminate mistakes.
Each team produced a set of annotation results. GPT-4-0125 annotated the same 13k (aspect, sentence) triplets. We calculated Exact Accuracy, F1-binary, Spearman Correlation, and Kappa Coefficient between the two human teams and between each human team and GPT-4. Table 2 shows human-human correlations in the low-to-mid 70s, indicating reasonable agreement, despite the fact that finding sentences that cover aspects is non-trivial due to different annotators’ potentially varied interpretations of key terms in Definition 1 such as "majority" and "reasonably inferred". For context on inter-annotator agreement in similar aspect-based tasks, Wang et al. reported 54% $\rho$ (three-way annotation) and Mysore et al. found near 40% pre-adjudication $\rho$ (four-way annotation).
Table 2 shows human-GPT4 and human-human correlations are comparable. A post-hoc analysis reveals that powerful embedding models like OpenAI-v3-large OpenAI (2024) have significantly lower correlation with humans. For the primary human-GPT4 comparison, the p-value (p = 0.3) from our bootstrapped test fails to reject the null hypothesis (no difference in correlation). In contrast, we can reject the null hypothesis for the embedding model which has clear and detectable differences from human annotators.
4.4 Support Sentence Retrieval (SSR)
With our annotation pipeline, we transform a support judgment into a special information retrieval task that evaluates if a model can identify and retrieve these support sentences from the full text of the paper. Since these support sentences form the basis of the support judgment, if a model can locate these support sentences, it is one step closer to reaching a good support judgment and classifying the correct risk-of-bias level.
Figure/table captions are split into sentences as text. Tables are included and turned into markdown format and treated as individual sentences. Figures are excluded since SSR is designed to evaluate textual models.
We have obtained the Aspect-to-Sentence Mapping. From this mapping, there exists the smallest optimal integer $K$ such that there are $K$ sentences that can cover all aspects in a support judgment. Therefore, when a model is only allowed to retrieve $K$ sentences from a study, we measure the percentage of aspects that the model’s retrieved set of sentences can cover. This metric is denoted as Aspect Recall Ratio @ Optimal.
Formally, let $\{s_{1},\dots,s_{k}\}$ be the set of retrieved sentences, and $\{f_{1},\dots,f_{m}\}$ be the set of aspects. We define the indicator function $\mathcal{S}(f_{i},s_{j})$ as:
$\mathcal{S}(f_{i},s_{j})=\begin{cases}1&\text{if }s_{j}\text{ covers }f_{i}\\ 0&\text{otherwise}\end{cases}$
The Aspect Recall Ratio @ Optimal is defined as
$\frac{\sum_{i=1}^{m}\displaystyle\mathbf{1}\Bigg\{{\left(\sum_{j=1}^{K}\mathcal{S}(f_{i},s_{j})\right)\geq 1}\Bigg\}}{m}$ (1)
Not only does it evaluate if a model can find support sentences, but it also evaluates if a model can control information redundancy from these sentences. Figure 5 in Appendix C visualizes SSR.
Bias Type
Model Avg Selection Attrition Performance Detection Reporting
n = 157 n = 46 n = 54 n = 61 n = 14
OpenAI-v3 22.7 40.1 6.3 26.7 27.4 13.1
GritLM-7B 18.9 35.8 7.0 15.2 26.8 9.5
GPT-4o 47.5 60.5 42.4 41.5 50.1 43.0
Sonnet-3.5 39.2 55.4 30.7 37.0 37.2 35.5
Llama-3.1-70B 45.6 61.2 34.2 45.2 50.1 37.4
Llama-3-8B 22.7 49.4 14.5 22.1 20.8 6.5
Llama-3-8B 40.8 69.0 24.8 47.3 48.6 14.3
Fine-tuned
Table 3: Support Sentence Retrieval (SSR). Evaluated on Cochrane Test (SSR). Metric is Aspect Recall Ratio @ Optimal. Llama-3-8B fine-tuned on Cochrane Train (SSR). Fine-tuning details are included in Appendix E.
Bias Type
Model Avg Selection Attrition Performance Detection Reporting
n = 130 n = 98 n = 87 n = 82 n = 80
GPT-4o 47.2 58.5 60.2 48.3 42.7 26.3
Sonnet-3.5 59.9 73.1 73.5 50.6 51.2 51.3
Llama-3.1-70B 53.2 66.2 62.2 44.8 46.3 46.3
Llama-3-8B 26.5 26.9 34.7 24.1 22.0 25.0
Llama-3-8B 29.6 40.8 28.6 24.1 19.5 35.0
Fine-tuned
Table 4: Support Judgment Selection (SJS). Evaluated on Cochrane Test (SJS). Metric is Accuracy. Llama-3-8B fine-tuned on Cochrane Train (SJS). Fine-tuning details are included in Appendix E.
Bias Type
Model Avg Selection Attrition Performance Detection Reporting Deviation
n = 933 n = 629 n = 309 n = 645 n = 594 n = 331
GPT-4o 42.1 50.5 36.5 54.7 43.7 34.4 32.7
Sonnet-3.5 41.9 52.8 37.1 50.6 43.3 33.6 34.2
Llama-3.1-70B 38.8 48.1 31.7 48.0 44.4 31.6 29.0
Llama-3-8B 30.1 36.4 32.4 39.1 37.2 19.8 15.4
Llama-3-8B Fine-tuned 36.3 49.5 33.3 42.4 40.1 27.0 25.7
LR Trained N/A 26.0 21.7 27.9 28.6 21.8 N/A
SVM Trained N/A 30.7 24.2 23.3 25.6 23.5 N/A
Table 5: Main Task, Risk-of-Bias Determination. Evaluated on Cochrane Test (Main) + Non-Cochrane Test (Main). Metric is Macro-F1. Llama-3-8B, logistic regression, and SVM are fine-tuned/trained on Cochrane Train (Main). Fine-tuning details are included in Appendix E.
4.5 Support Judgment Selection (SJS)
Simply retrieving support sentences is insufficient for a risk-of-bias determination. A model needs the ability to interpret and synthesize information from these retrieved support sentences in order to generate a good support judgment that will lead to the correct determination of risk-of-bias.
Following the convention of using multiple-choice questions as a proxy for generative tasks Rein et al. (2023), we propose the support judgment selection subtask, an MCQ task, where the model selects the correct support judgment from a mix of three synthetically generated support judgments, and three human-written options derived from other papers’ support judgments for the same exact bias. These three human-written options sometimes proved to be surprisingly hard distractors because their lack of direct paper quotes made them into generic, non-paper-specific statements. A deep logical understanding of the paper’s content rather than superficial semantic matching is required to look past them and find the correct one.
We prompt GPT-4 to generate the three synthetic options by imitating support judgments from other papers concerning the same exact bias name. These options are tailored to be paper-specific while maintaining the underlying reasoning. Empirically, these six distractor options are often misleading. Table 4 shows the best LLM only achieves 60% accuracy. In order to ensure that all incorrect options were actually incorrect, all 906 datapoints were checked by human, and 465 datapoints were kept. See Figure 6 in Appendix C to visualize SJS.
@@ 5 Experiments
5.1 Evaluated Models
We evaluate various LLMs, GPT-4o-2024-05-13 OpenAI (2024a), Sonnet3.5-20240620 Anthropic (2024), and Llama-3.1-70B Dubey et al. (2024). We evaluate and fine-tune light-weight Llama-3-8B Dubey et al. (2024). We also evaluate two recent embedding models GritLM 7B Muennighoff et al. (2024) and OpenAI-embedding-v3-large OpenAI (2024). Note the GPT-4o model is different from the GPT-4-0125 model used in our annotation pipeline.
For the main task, we further evaluate three traditional ML models: RobotReviewer Marshall et al. (2016), Support Vector Machine (SVM) and Logistic Regression (LR) Dias et al. (2025).
For all generative language models, we use chain-of-thought as a prompting strategy to stabilize the model performance. We develop our prompts on train sets. For reproducibility, we use a common generation setting of temperature 0, top-p 1, frequency penalty 0, and presence penalty 0.
@@ 5.2 Evaluation Metrics
Some bias names may fall into multiple categories, causing the total datapoint count across categories to be slightly higher than the test set size. Model performance is reported per category and as an average across all categories.
The SSR subtask uses Aspect Recall Ratio @ Optimal defined in Equation 1 as its metric. It measures the percentage of aspects covered by the retrieved sentences. Unlike traditional metrics (e.g., recall, precision), it penalizes redundancy by rewarding models that retrieve the minimum number of sentences (typically 1-2) needed to cover all aspects for a given data point. For instance, if one sentence covers multiple aspects, retrieving additional sentences that cover already addressed aspects is discouraged.
6 Benchmarking Results and Analyses
6.1 LLMs Exhibit Distinct Strengths in Retrieval and Reasoning
While GPT-4o and Sonnet-3.5, with 42.1% and 41.9% average Macro-F1, are comparable in the main task Table 5, a closer inspection of their performances on the two subtask Tables 4, 4 reveals interesting differences. In support sentence retrieval (SSR), GPT-4o has 47.5% Aspect Recall Ratio, while Sonnet-3.5 only has 39.2%. This shows that Sonnet-3.5 is worse at retrieving the best support sentences that are the bases of human expert-level support judgments. However, in support judgment selection (SJS), Sonnet-3.5 outperforms other GPT-4o with 59.9% over 47.2% accuracy. This shows that Sonnet-3.5 can select better support judgments, which involves reasoning and deliberating which option best aligns with the paper’s contents and the bias and risk-level definitions.
In other words, GPT-4o’s advantage over Sonnet-3.5 is information retrieval, while Sonnet-3.5’s advantage is information synthesis with reasoning. We provide an example in Appendix G comparing reasoning traces of the two models to clearly illustrate Sonnet-3.5’s advantage in reasoning.
6.2 Positive Correlation Between Main Task and Subtasks
We investigate the relationship between the main task and our subtasks. For an LLM (either GPT-4o or Sonnet 3.5), we categorize datapoints in the SSR subtask into two classes: datapoints whose risk-of-bias level is correctly predicted by the LLM and those incorrectly predicted. For each of the two classes, we calculate its average Aspect Recall Ratio @ Optimal. See the left part in Figure 4, we observe that the Aspect Recall Ratio is significantly higher in the correct class than the incorrect class, for both Sonnet-3.5 and GPT-4o, suggesting LLMs’ performances on SSR are positively correlated with their performances on the main task.
Similarly, datapoints in the SJS subtask are divided into two classes. From the right part in Figure 4, we see that, for both Sonnet-3.5 and GPT-4o, their performances on SJS are positively correlated with their performances on the main task.
Figure 4: There is a positive correlation between the main task and SSR & SJS subtask performance.
6.3 Retrieval and Reasoning Ablation
Retrieval Reasoning Main Task Result
Sonnet-3.5 GPT-4o 31.5 $\pm$ 5.0
GPT-4o GPT-4o 40.3 $\pm$ 6.9
Ground Truth GPT-4o 46.1 $\pm$ 7.5
Ground Truth Sonnet-3.5 49.4 $\pm$ 8.0
Table 6: Two stage pipeline that first retrieves and then reason based on the retrieved results (full paper is not shown). Main Task, Risk-of-Bias Determination. Evaluated on the 313 datapoints in Cochrane Test (SSR), all of which have risk-of-bias labels. Metric is Macro F1 score (%) averaged across bias categories. The error bar represents the 95% confidence interval.
To investigate how the retrieval and reasoning abilities of LLMs affect their performance on the main task, we employ a two-stage retrieval and reasoning pipeline. In this setup, one LLM processes sentences retrieved by another LLM, rather than accessing the original paper directly. The specific prompt used is detailed in Appendix I.3. This experiment was conducted on Cochrane Test (SSR), which comprises 313 datapoints, each associated with a support judgment and a risk-of-bias label.
We pair different models’ retrieved sentences with different reasoning models to evaluate performance on the main task Risk-of-Bias Determination. Ground truth retrieved sentences are provided by our annotation pipeline.
As shown in Table 6, when results from different retrieval methods are paired with the same reasoning model (GPT-4o), ground-truth retrieval achieves significantly higher macro F1 score (46.1%) over Sonnet-3.5’s retrieval (31.5%), confirming that retrieval quality impacts main task performance. When the retrieval quality is held constant (using ground-truth retrieval) and the reasoning model varied, there is some inconclusive evidence that suggests Sonnet-3.5 slightly outperforms GPT-4o. These ablations demonstrate that both retrieval and reasoning abilities are relevant for the Risk-of-Bias Determination task.
6.4 Retrieval Marginally Improves LLM on Non-Cochrane Test Set
So far we have focused on the Cochrane test set, which includes support judgments, unlike the Non-Cochrane test set. To investigate whether more accurate support sentences could enhance LLM performance on the Non-Cochrane test set (2489 datapoints), we used the o4-mini reasoning model OpenAI (2024b) to find such sentences within each paper, guided by the datapoint’s actual risk level (see Appendix I.4 for the prompt). In this proof-of-concept setup, the LLM received the full paper and o4-mini-identified support sentences as highlights, which we compared against baselines with full paper but no highlights.
Table 7 reveals that both GPT-4o and Sonnet-3.5 showed slight performance improvements on the Non-Cochrane Test (Main) over their baselines. The limited gain is partly due to inaccuracies in retrieved sentences, as recovering support sentences without expert support judgments is challenging.
Setting Main Task
GPT-4o Full Paper 42.5 $\pm$ 2.2
GPT-4o Full Paper + Highlight Retrieval 44.1 $\pm$ 2.2
Sonnet-3.5 Full Paper 40.3 $\pm$ 2.2
Sonnet-3.5 Full Paper + Highlight Retrieval 44.5 $\pm$ 2.2
Table 7: Main Task, Risk-of-Bias Determination. Evaluated on Non-Cochrane Test (Main). Metric is Macro F1 score (%) averaged across bias categories. The error bar represents the 95% confidence interval.
6.5 Challenges in Support Sentence Retrieval
Our findings establish the retrieval task SSR as a crucial intermediate milestone for improving LLM on the main task. Nevertheless, Table 4 indicates that in zero-shot setting, current models struggle significantly: modern embedding models such as OpenAI-v3 and GritLM-7B achieve low recall ($\leq$22.7%), while commercial LLMs also underperform (recall $\leq$47.5%). Although fine-tuning LLama3-8B boosts its SSR recall from 22.7% to 40.8%, its retrieval quality remains unsatisfactory.
A significant limitation of embedding models is their lack of context awareness. Even recent contextual-aware models Morris and Rush (2024) cannot handle the long-range interactions across different sentences. In Appendix F, we examine a type of globally-connected datapoints in SSR potentially necessitates a more global understanding of the entire context. We leave the exploration of contextual embedding models and more accurate long-context LLMs for solving SSR as a future research direction.
6.6 Limitations of Traditional ML Models on RoBBR
Biases
Model Avg Allocate Blind Blind RanSeq
Conceal Outc Part Gen
n=32 n=19 n=18 n=30
RobotReviewer 56.7 $\pm$ 8.4 75.0 39.1 43.8 68.9
LR 53.1 $\pm$ 9.7 71.9 50.4 51.8 38.4
SVM 44.8 $\pm$ 8.6 45.9 55.2 41.9 36.0
GPT-4o 65.6 $\pm$ 8.5 83.6 59.1 41.9 77.8
Sonnet-3.5 67.5 $\pm$ 8.4 77.0 82.5 41.9 68.8
Table 8: Main Task, Risk-of-Bias Determination. Evaluated on a subset from the Cochrane Test (Main), comprising only the four RobotReviewer-assessable bias types: allocation concealment, blinding of outcome, blinding of participants, and random sequence generation. Only two judgment categories (low, and high/unclear) per RobotReviewer’s specification. Metric is Macro-F1. The error bar represents the 95% confidence interval.
This study compares previous ML models with LLMs on four specific biases, the only ones assessable by RobotReviewer Marshall et al. (2017) (an SVM/CNN ensemble) due to its inaccessible training code. We also trained logistic regression and SVM models on our Cochrane Train Set (Main) Dias et al. (2025). Table 8 shows inconclusive results : RobotReviewer, SVM, and LR perform marginally worse, if at all, than GPT-4o and Sonnet-3.5.
The four biases RobotReviewer assesses are considered straightforward because superficial keywords often directly indicate bias risk (e.g., "opaque envelope" for low allocation concealment bias; "randomness" or "random number generator" for low random sequence generation bias), suggesting significant keyword reliance. However, biases like selective reporting or deviation bias may necessitate a more holistic consideration of the entire biomedical paper.
Traditional ML models by definition are hard to generalize to new bias names and definitions. RoB1/RoB2 guidelines list hundreds of bias names, some with definitions re-interpreted in different reviews. we treat these as paper-specific biases, manually incorporating the paper’s re-interpreted definition during LLM evaluation on RoBBR. Thus, RoBBR primarily evaluates models in zero-shot settings or fine-tunes models like Llama3-8B for generalization to unseen bias definitions via semantic understanding. Furthermore, most bert-based models’ 512-token context limit is ill-suited for RoBBR’s biomedical studies (average 8k tokens), unlike modern LLMs with longer context windows.